1. Two wagons, with masses 40t and 50t, move in opposite directions with speeds respectively 7m / s and 5m / s. After hitting the wagons continue to move together. Find the size and direction of movement of the wagons after the impact. Friction is not taken into account.


1
Expert's answer
2021-02-11T10:41:11-0500

Let's choose the right as the positive direction. Let the 40t wagon moves to the right and the 50t wagon moves to the left. Then, from the law of conservation of momentum, we have:


"m_1v_1-m_2v_2=(m_1+m_2)v,""v=\dfrac{m_1v_1-m_2v_2}{m_1+m_2},""v=\dfrac{40000\ kg\cdot7\ \dfrac{m}{s}-50000\ kg\cdot5\ \dfrac{m}{s}}{40000\ kg+50000\ kg}=0.33\ \dfrac{m}{s}."

The sign plus means that the wagons will move to the right after the impact.